OpenStudy (anonymous):

What is log2(x-6) + log 2(9)=2 What is x?

OpenStudy (anonymous):

log2 [ 9(x-6) ] = 2

OpenStudy (anonymous):

9(x-6) = 2^2

OpenStudy (anonymous):

\[The 2 next \to \log is the opposite of exponent. Its a little 2 under the \log. Dont know how \to type that...\]

OpenStudy (anonymous):

x=58/9

OpenStudy (anonymous):

Log (x-6) + log (9) = 2 2 2

OpenStudy (anonymous):

How about log 2 + log (x-10)=1 This one is written just as I typed. No little opposite of exponents

OpenStudy (anonymous):

log [2(x-10)] =1 log [2x-20] =1 Assuming its base 10 , which is what is usually assumed if there is no base

OpenStudy (anonymous):

Change of base I mean

OpenStudy (anonymous):

2x-20 =10^1 2x = 30 x=15

OpenStudy (anonymous):

Thank you